OpenStudy (anonymous):

what is (6/(x)^-1+2)

OpenStudy (anonymous):

Do you mean,\[\frac{6}{x^{-1}}+2\]?

OpenStudy (anonymous):

the +2 is in the denominator

OpenStudy (anonymous):

(6/(x)^-1+2) = 6/[(1/x)+2]=6x/(1+2x)

OpenStudy (anonymous):

\[\frac{6}{x^{-1}+2}=\frac{6}{\frac{1}{x}+2}=\frac{6}{\frac{1+2x}{x}}=6 \times \frac{x}{1+2x}=\frac{6x}{1+2x}\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Fan me :D